The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs.
I am so stuck I do not even know where to start. Please help!!

Answer by tutorcecilia(2152)

Sketch the triangle and lable the parts.
.
.
P=Side #1 + Side #2 + Side #3 [Formula for the perimeter of a triangle]
Since it is an isoceles triangle, two of the sides will be equal.
Perimeter=29
Side #1 = (4 x base)-8=4b-8
Side #2=(4 x base)-8=4b-8
Side #3=the base=b
.
P=(4b-8) + (4b-8)+(b)
29=4b-8 + 4b - 8 + b
29=8b-16+b
29=9b-16
29+16=9b-16+16
29+16=9b
45=9b
45/9=b
5=b
.
Plug (b=5) back into (4b-8) to find the length of the equal sides:
4b-8
4(5)-(8)
20-8
12= length of one leg
.
Check by plugging all of the values back into the formula for the perimeter of a triangle:
P=s+s+s
P(4b-8)+(4b-8)+b
29=(4)5)-8+(4)(5)-8+5
29=20-8+20-8+5
29=12+12+5
29=29 [Checks out]